Design a three-mode temperature control system which inputs error

in 0-4 V range. The output to final control element is 0-8 V. Given:


Kp = 2.4 % per %

Ki = 9 % (%/min)/ %


5


Kd = 0.7 % / (%/min)


1
Expert's answer
2021-05-31T06:40:12-0400

kp"= 2.4" % will be "\frac{100}{x}" % "=x=\frac{100}{2.4}=41.6" %

"GP=" "\frac{(100\%)(V_{out(max)}-V_{out(min)}}{(41.6\%)(V_{in(max)-V_{in(min)}}}=\frac{(100\%)(8V)}{(41.6\%)(4V)}=4.8"

"=K_D=0.7" %

"[\frac{0.7}{\%min}][\frac{60sec}{1min]}=42\frac{\%}{\%/sec}"

GD"=\frac{(42\%)8v}{1\frac{\%}{sec}(4v)}"

Now period of fastest expected change"=8sec"

"\therefore \frac{R_1}{R_1+R_3}=R_3C=\frac{0.1}{2\pi}\times 8sec=0.1273sec"

We have the relation , GP"=[\frac{R_2}{R_1+R_3}]=4.8"

"G_D=R_3C=60S ec"

"\frac{R_1}{R_1+R_3}=1\%4V=0.004" and "C=10\mu F"

We will have "R_3=6m\Omega \therefore R_3C=60 sec \space and R_3=\frac{60}{10}=6"

"R_1=0.004R_1+0.004R_3"

"R_1=0.004R_1=0.004(6\Omega)"

"0.996R_1=24k"

"R_1=\frac{24}{0.996}=24.09k\Omega"

"\frac{R_2}{R_1+R_3}=4.8"

"R_2=4.8(R_1)+4.8R_3"

"=4.8(24.09)+4.8(6)"

"=144.432m\Omega"